OpenStudy (anonymous):

State which set of three numbers could represent the lengths of the sides of a right triangle. A.) 8, 12 , 16 B.) 16, 32, 36 C.) 5, 12, 13 D.) 6, 7, 8

HanAkoSolo (jamierox4ev3r):

@lisa2005 let's keep in mind the pythagorean rule: a^2 + b^2 = c^2 if any of the numbers fit this description, it could be a right triangle.

HanAkoSolo (jamierox4ev3r):

For example, let's take the numbers of D. 6,7,8, plug them into the pythagorean theorem 6^2 + 7^2 = 8^2 which becomes 36 + 49 =64 Is this a true statement? unfortunately, no, so D is not correct

HanAkoSolo (jamierox4ev3r):

Do what i did with D. Whichever one forms a true statement is correct :)

HanAkoSolo (jamierox4ev3r):

@lisa2005 do you understand?

TheSmartOne (thesmartone):

Remember this rule: The sum of any 2 sides of a triangle must be greater than the third side.  Then \(\sf \Large a + b > c\\\Large a + c > b\\\Large b+ c > a\)

TheSmartOne (thesmartone):

@Jamierox4ev3r 's way will not work all the time, unfortunately. :P

TheSmartOne (thesmartone):

However, my way will always work :D

HanAkoSolo (jamierox4ev3r):

it works all the time if you're looking for a right triangle >_>

TheSmartOne (thesmartone):

Questions like these however are not limited to right triangles all the time. :P